An empty pressure cooker of volume 10 littee contains air at atmospheric pressure 10^5 pascal and temp of 27°C.It contains a whistle which has area of 0.1 cm^2 and weight of 100 gn .what should be temp of air inside so that the whistle is just lifted up. 

Grade:12th pass

1 Answers

Arun
4 years ago
 
Pressure by whistle = Po + F / A = Po​ + mg/ A
= 105Pa + (0.1 * 10 N) / (0.1 * 104 m2)
= 2 * 105 Pa
Pressure inside cooker = 105Pa
for whistle to just lift, by gay lussacs law
Pwhistle / Tsurrounding = Pcooker / Tcooker
(2 * 105 Pa) / 300 K = 105 Pa / T
T = 600 K = 327oC